# Governing Together: Toward Infrastructure for Community-Run Social Media

### v1

@@ 8. Conclusion
Drawing from four workshop sessions with Fediverse organizers and developers, we make a case for inter-community governance in social computing systems, presenting a vision of concrete ideas and principles that may guide future design. Community-level decision-making plays a critical role in enacting governance in social computing systems, but faces persistent challenges and harms. Inter-community governance offers a ‘meta’ approach to addressing these issues by envisioning an infrastructural layer that support interaction and relationships among communities, so that they may govern together — even while autonomous. Focused on the case of community-run social media, discussions with participants underscore six challenges that inter-community governance grapples with: making community governance decisions visible; collating nuanced information about issues; managing what information shared, to whom; enabling new relationships between communities; being able to customize community governance decisions; and minimizing technical barriers to adoption.
To meet these challenges in design, we suggest that future design efforts focus on three interdependent principles: modularity, forkability, and polycentricity. By synthesizing the ideas discussed across workshops into an ecosystem of concrete tools, interfaces, and mechanisms, we demonstrate one way these principles might operate together. The imagined ecosystem maps a path forward for future work to interrogate and experiment with inter-community governance, toward building infrastructure that supports communities in a variety of social computing systems.
@@ Acknowledgements.
We are grateful to our participants who were generous with their time and thoughts in workshops, follow-up calls, and beyond.

### v2

@@ 8. Conclusion
Drawing from four workshop sessions with Fediverse organizers and developers, we make a case for inter-community governance in social computing systems, presenting a vision of concrete ideas and principles that may guide future design. Community-level decision-making plays a critical role in enacting governance in social computing systems, but faces persistent challenges and harms. Inter-community governance offers a ‘meta’ approach to addressing these issues by envisioning an infrastructural layer that support interaction and relationships among communities, so that they may govern together while preserving autonomy. Focused on the case of community-run social media, discussions with participants underscore six challenges that inter-community governance grapples with: making community governance decisions visible; collating nuanced information about issues; managing what information shared, to whom; enabling new relationships between communities; being able to customize community governance decisions; and minimizing technical barriers to adoption.
To meet these challenges, we suggest that future inter-community governance design efforts focus on three interdependent principles: modularity, forkability, and polycentricity. By synthesizing the ideas from workshops into an ecosystem of tools, interfaces, and mechanisms, we demonstrate one way these principles might operate together. The imagined ecosystem maps a path forward for future work to interrogate and experiment with inter-community governance, toward building infrastructure that supports communities in a variety of social computing systems.
@@ Acknowledgements.
We are grateful to our participants who were generous with their time and thoughts in workshops, follow-up calls, and beyond. We thank our thoughtful reviewers for their constructive feedback on this piece. An in-progress version of this work was shared at a Siegel Research Fellow meeting as part of the first author’s fellowship.
Appendix A Workshop Protocol
The workshop began with a warm-up example of cooperation in commons-based resource management based on the work of Acheson (1988), presented with the slide shown in Figure 1. We asked: Go around and share anything loose thoughts about what we noticed, wondered about, or found interesting with this example. What made cooperation possible, and what were the challenges they faced in their context?
Next, we introduced the concept of governance frictions, which we defined as “refer to incompatibilities in community governance decisions that enforce the goals of one community but undermine the goals of another.” The workshop then moved into a series of 20-30 minutes of breakout discussions, with the following written prompts:
  1. (1)
    Governance frictions seen and encountered in the wild: What are some concrete, specific decisions made about how to run and govern a community on the Fediverse? These can be social, technical, organizational, etc.
    • Narrow down to 2-3 decisions you feel are the most pressing and important. Why?
    • For each decision: what choices might other communities make about that decision? What problems or challenges result, if at all? Who deals with them?
  2. (2)
    Ideal workflows: Create a flow chart of steps in an indeal workflow for handling the friction your group has selected. You’re in an ideal world: perhaps your workflow is part of new policies, technologies, agreements, infrastructures, etc. Break down into the steps and per step, decisiosn that need to be made.
  3. (3)
    Tools to support our ideal workflow: (1) Brainstorm a speculative tool that supports our ideal workflow in some way. (2) Tell us how it works, including: the goal (+what step/s of the workflow it’s a part of), features, limitations, who is using the tool, who makes decisions about it, any relevant policies or processes. Remember we’re in an ideal world - the tool might not be immediately possible in our world!
We verbally noted to participants that by focusing on ideal workflows and speculative tools, we aimed to imagine possible sociotechnical futures - not only focus on how governance currently works on the Fediverse. For Workshops 3 and 4, Activity 2 and 3 were merged into one activity, as we found that participants tended to talk about speculative tools immediately with workflows, and for time.
Appendix B Summary of Ideas from Workshops
The following tables provide a succinct summary of core points of discussion in the workshops, particularly the kinds of decisions participants felt were important and concepts/potential tools or resources ideated on.
Technical and Infrastructural Decisions use of complementary platforms like Loomio
server and hosting infrastructure
what relays used
tools like bots
platform software
member-facing features
Organizational Decisions number of admins
ownership
internal communication norms
formalization as an entity
funding model
legal obligations
governance capacity
decision-making model (e.g., cooperative)
rule enforcement process
Normative Decisions codes of conducts, rules, and the concreteness of those rules
content warning requirements
terms of service
use of blocklists
documentation of community practices (for transparency)
membership (e.g., open vs closed registration)
onboarding practices
federation policy
bridging (to other services)
open to threads/meta, mastodon.social, and other at-scale services
relationship to third-party groups like IFTAS
privacy policy
data permanence and maintenance
Table 2. A summary of community decisions listed across the four workshops. Participants noted that a community’s specific goals shaped all of these decisions.
Groupings of community decisions mentioned in the workshops. The decisions are in three groupings: technical and infrastructural decisions; organizational decisions; and normative decisions. Technical infrastructural decisions contain things like: relay, server and other hosting infrastructure, platform software, tools like bots, member-facing options. Organizational decisions contain things like: number of admins, ownership, legal obligations, decision-making model and processes. Normative decisions have things like: federation policy, terms of service, bridging, blocklist, data permanence and maintenance.
Concept W1 W2 W3 W4
Nutrition Facts shared vocabulary of governance encoded, standardized code of conduct; forking; mapping compatibility across communities guide of community governance with a nutrition label
Trust Bubbles RSS feed for a circle of communities an “archipelago” of multiple coalitions; local bubbles and “community of communities” sharing moderation tasks within a bubble, ideally integrated into existing software “ladder of engagement” based on a “chain of trust” that de-dichotomizes the concept of federation
Shared Logs moderation and blocking decisions in a public log archive, possibly represented as abstractions of patterns that need review mechanism for pooling moderation tasks moderator dashboard briefly mentioned
Issue Bulletin meta-community; shared console, using algorithmic sorting to highlight trending and relevant discussions a space for bouncing ideas and having discussions about issues mod-mod communication tools, like a moderator inbox sharing, verifying, collating, and triaging complaints and evidence of harm
Table 3. High-level summary of how concepts from the ecosystem relate to more substantive discussions across the workshops (W#), briefly noting the more specific or particular ways participants talked about the concept. Note that not all concepts were explored equally in depth in every workshop, and all of the concepts are somewhat tied to each other.